# MOFFlow: Flow Matching for Structure Prediction of Metal-Organic Frameworks

### v1

@@ 5.2 Property evaluation
For comparison, we use DiffCSP as a representative of general CSP approaches. We exclude optimization-based baselines as they did not yield meaningful results. For each test structure from Section 5.1, we generate a single sample and compute the relevant properties using Zeo++ (Willems et al., 2012). We do not filter our samples with the match criteria from Section 5.1. We use 50 integration steps for our model and measure performance with RMSE.
@@ 5.4 Comparison to self-assembly algorithm
Results. Table 3 shows that MOFFlow alone outperforms the combination with the self-assembly algorithm, indicating that learning the building block orientations leads to more accurate MOF structure predictions than heuristic-based overlap optimization. Additionally, our method offers faster inference, further demonstrating its efficiency compared to optimization-based approaches.
6 Conclusion
@@
We describe experimental details and hyperparameters in Appendix C. We provide our codes and model checkpoint in https://anonymous.4open.science/r/MOFFlow-3547.
Ethic Statement
Our framework targets to advance porous material discovery, which is deeply related to carbon capturing, catalysis design, and drug discovery. We believe our work can improve the quality of human life by assisting in resolving global warming and drug design. However, they should notice the caution due to the misuse in developing hazardous materials or products that may be harmful for specific usage.
@@ Appendix C Implementation details
Training details. We employ the $\operatorname{TimestepBatch}$ algorithm (Yim et al., 2023b) to efficiently construct batches. This method generates a batch by applying multiple noise levels $t\in[0,1]$ to a single data instance, ensuring uniform batch size. To manage memory constraints, we cap the batch size with $N^{2}$, where $N$ is the number of atoms. The model is trained on eight 24GB NVIDIA RTX 3090 GPUs over 15 epochs, taking approximately 7 days.
@@
Baselines. For DiffCSP (Jiao et al., 2024a), we change the edge construction method from original fully connected to a radius graph with a cutoff of 5Å since a fully connected graph is not feasible for large-size crystals like MOFs. We use batch size 8 to cope with memory constraints. Any other options followed the default hyperparameter of the DiffCSP. We trained 200 epochs, taking approximately 7 days until convergence on a 24GB NVIDIA RTX 3090 GPU.
For random search (RS) and evolutionary algorithms (EA) for structure prediction, we considered CrySPY (Yamashita et al., 2021) software and deep-learning based structure optimizer (Deng et al., 2023). We set the symmetry-based search for RS and constructed 20 structures per sample. Also, we set the EA for 5 initial RS and ran a maximum of 20 generations with 10 crossovers, 4 permutations, 2 strains, and 2 elites for 20 populations. We considered the tournament-selection function for 4 tournament sizes. We used energy-based optimization for RS and EA, since there exists no publicly opened software for MOF crystal structure prediction.
Code. Our implementation is built upon https://github.com/microsoft/protein-frame-flow, https://github.com/microsoft/MOFDiff, https://github.com/gcorso/DiffDock, and https://github.com/vgsatorras/egnn. We appreciate the authors (Yim et al., 2023a; b; 2024; Fu et al., 2023; Corso et al., 2023; Satorras et al., 2021) for their contributions.
@@
Appendix D Defining local coordinates of building blocks
@@
Appendix E Model architecture

### v2

@@ 5.2 Property evaluation
We compare our model to DiffCSP as a representative general CSP approach. We exclude optimization-based baselines as they did not yield meaningful results. For each test structure in Section 5.1, we generate a single sample with 50 integration steps, then evaluate the properties of predicted and ground-truth structures using Zeo++ (Willems et al., 2012). We evaluate the models with RMSE and distributional differences. We do not filter our samples with the match criteria from Section 5.1.
@@ 5.4 Comparison to self-assembly algorithm
Results. Table 3 shows that MOFFlow alone outperforms the combination with the self-assembly algorithm, indicating that learning the building block orientations leads to more accurate MOF structure predictions than heuristic-based overlap optimization. Additionally, our method offers faster inference, further demonstrating its efficiency compared to optimization-based approaches. A comprehensive comparison is provided in Appendix H.
6 Discussions
@@
Limitations. MOFFlow was evaluated only on the hypothetical database (Boyd et al., 2019), highlighting the need for evaluation on more challenging real-world datasets. Additionally, we assume that the local building block structures are known (i.e., the rigid body assumption), which may be impractical for rare building blocks whose structural information is missing from existing libraries (e.g., Gibaldi et al. (2022)) or is inaccurate. Finally, MOFFlow is not invariant to periodic transformations of the input; explicitly modeling periodic invariance could further improve its performance.
@@
We describe experimental details and hyperparameters in Appendix C. We provide our codes and model checkpoint in https://github.com/nayoung10/mofflow.
Acknowledgments
This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST)), National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2022-NR072184), and GRDC(Global Research Development Center) Cooperative Hub Program through the National Research Foundation of Korea(NRF) grant funded by the Ministry of Science and ICT(MSIT) (No. RS-2024-00436165).
@@ Appendix C Implementation details
Training details. We use the $\operatorname{TimestepBatch}$ algorithm (Yim et al., 2023b) to simplify batch construction. This method generates a batch by applying multiple noise levels $t\in[0,1]$ to a single data instance, ensuring uniform batch size. To manage memory constraints, we cap the batch size with $N^{2}$, where $N$ is the number of atoms.
@@
Baselines. Here, we provide the implementation details of the baselines. Unless specified otherwise, all hyperparameters follow their default settings.
  • DiffCSP (Jiao et al., 2024a): To address memory constraints, we replaced fully connected edge construction with a radius graph (cutoff: 5Å) and used a batch size of 8. The model was trained on a 24GB NVIDIA RTX 3090 GPU for 5 days until convergence.
  • RS & EA: Both random search (RS) and the evolutionary algorithm (EA) were implemented with CrySPY (Yamashita et al., 2021) and energy-based optimization with CHGNet (Deng et al., 2023). For RS, we generated 20 structures per sample with a symmetry-based search. EA began with 5 initial RS runs and performed up to 20 generations with a population size of 20, 10 crossovers, 4 permutations, 2 strains, and 2 elites. A tournament selection function with a size of 4 was employed.
FlowMM (Miller et al., 2024) and EquiCSP (Lin et al., 2024) were excluded from our baselines due to their high resource demands when adapted for MOF structure prediction, requiring more than 20 days of training on an 80GB A100 GPU.
Computational resources. Table 9 summarizes the computational resources required to train learning-based models. Notably, the $\operatorname{TimestepBatch}$ implementation of MOFFlow requires longer training times than DiffCSP in terms of GPU hours. To address this inefficiency, we also release a refactored $\operatorname{Batch}$ version of MOFFlow with details in Appendix D.
Codebase. Our implementation is built on https://github.com/gcorso/DiffDock, https://github.com/vgsatorras/egnn, https://github.com/microsoft/MOFDiff, and https://github.com/microsoft/protein-frame-flow. We appreciate the authors (Yim et al., 2023a; b; 2024; Fu et al., 2023; Corso et al., 2023; Satorras et al., 2021) for their contributions.
@@
Table 9: Computational resources. Comparison of batch size, GPU type ($\times$ number), GPU memory utilization (GB), and training time (d: days, h: hours) for training learning-based models.
Batch size GPU Type GPU memory (GB) Training time
DiffCSP 8 24GB 3090 (×$\times 1$) 23.24 - 23.96 5d 2h
MOFFlow ($\operatorname{TimestepBatch}$) 160 24GB 3090 (×$\times 8$) 6.67 - 23.95 5d 15h
MOFFlow ($\operatorname{Batch}$) 160 24GB 3090 (×$\times 8$) 21.62 - 23.80 1d 17h
Appendix D Batch implementation of MOFFlow
While the $\operatorname{TimestepBatch}$ algorithm (Yim et al., 2023b) simplifies implementation, it slows convergence due to its effective batch size of 1 (i.e., each batch contains noise-perturbed variants of a single instance). To address this limitation, we introduce the $\operatorname{Batch}$ implementation, which processes multiple data instances per batch, aligning with standard practice. As shown in Table 10, $\operatorname{Batch}$ achieves comparable performance to $\operatorname{TimestepBatch}$ while significantly reducing training time from $1087.25$ to $332.74$ GPU hours. It also reduces inference time from $1.94$ to $0.1932$ seconds.
Table 10: Comparison of $\operatorname{TimestepBatch}$ and $\operatorname{Batch}$ implementations of MOFFlow. The $\operatorname{Batch}$ implementation achieves comparable performance while significantly reducing training time (GPU hours) and inference time (seconds). Inference time is reported as the average per test instance, calculated by dividing the total elapsed time by the size of the test set.
Training time (h) Inference time (s) Match rate (%) RMSE
$\operatorname{TimestepBatch}$ 1087.25 1.94 31.69 0.2820
$\operatorname{Batch}$ 332.74 0.19 32.73 0.2743
Appendix E Defining local coordinates of building blocks
@@
Appendix F Model architecture
@@
Appendix G Effect of integration steps
Here, we investigate how the number of sampling integration steps affects MOFFlow’s performance. We randomly select 1,000 structures from the test set used in Section 5.1 and evaluate the match rate, RMSE, and average sampling time for varying integration steps: $(5,10,50,100,200)$. For each structure, We generate a single sample and set the thresholds for $\operatorname{stol}$, $\operatorname{ltol}$, and $\operatorname{angle\_tol}$ to $(0.5,0.3,10.0)$.
Results. Figure 7 presents our results. Notably, the performance peaks around $10$ and $50$ integration steps, with a slight decline observed for higher step counts. This aligns with the trends reported by Miller et al. (2024). Based on these results, we use $50$ integration steps in our main experiments, which yield the highest match rate of $31.1\%$, a low RMSE of $0.2821$, and a fast sampling time of $1.267$ seconds.
Figure 7: Effect of integration steps on match rate, RMSE, and average sampling time. Performance is highest at $10$ and $50$ integration steps. We select $50$ steps for the main experiments due to its optimal balance of the highest match rate, low RMSE, and efficient sampling time.
Appendix H Comparison with self-assembly algorithm
We compare MOFFlow’s scalability and sampling efficiency with the self-assembly algorithm (Fu et al., 2023), highlighting their respective strengths and limitations.
Scalability. Since both methods operate at the building block level, we compare their match rates as a function of the number of building blocks. The experimental settings follow Section 5.4. Figure 8(a) shows that while MOFFlow achieves a higher overall match rate ($31.69\%$ vs. $27.14\%$), the self-assembly algorithm scales better for structures with more building blocks.
Sampling efficiency. Figure 8(b) compares the assembly times of the two methods. MOFFlow (left) demonstrates significantly faster sampling, below 8 seconds per sample. In contrast, the self-assembly algorithm (right) often requires over 2000 seconds per sample. Additionally, the self-assembly algorithm’s performance varies widely with initialization, with average assembly times ranging from 4.75 to 14.62 seconds per trial, reflecting its sensitivity to initial conditions.
(a) Match rate by building block count
(b) Distribution of assembly time
Figure 8: Comparison of scalability and efficiency between MOFFlow and the self-assembly algorithm. (8(a)) Match rate across varying building block counts. MOFFlow achieves higher match rates overall, but the self-assembly algorithm performs better for structures with a large number of building blocks. (8(b)) Assembly time distributions for (left) MOFFlow and (right) the self-assembly algorithm, highlighting the significantly faster inference speed of MOFFlow.